Assume the region to the right of a certain plane
Answer #1

given,

magnetic field = 1.98 mT

time interval = pi * mass of electron / charge of electron * magnetic field

time interval = pi * 9.109 * 10^-31 / (1.6 * 10^-19 * 1.98 * 10^-3)

time interval = 9.03307054 * 10^-9 sec

r = 1.97 cm

r = mv / qB

0.0197 =  9.109 * 10^-31 * v / (1.6 * 10^-19 * 1.98 * 10^-3)

v = 6.851 * 10^6 m/s

kinetic energy = 0.5 * mv^2

kinetic energy = 0.5 * 9.109 * 10^-31 * (6.851 * 10^6)^2

kinetic energy = 2.1377 × 10^-17 J